OpenStudy (anonymous):

The roots of the equation k^4 - 7k^2 - 8=0 are?

OpenStudy (callisto):

k^4 - 7k^2 - 8=0 (k^2)^2 -8k^2 + k^2 -8 = 0 k^2 (k^2 -8) + (k^2 -8) = 0 (k^2 +1) ( k^2 -8) = 0 k^2 +1 = 0 or k^2 -8 = 0 Now solve these two equations.

OpenStudy (anonymous):

thxxx for helping i got it

OpenStudy (callisto):

welcome!